antitative Problem: Bellinger Industries s considering two ro ects or inclusion in its capital budget and you have been asked do the anal sis th proje a ter tax cas no s are shown on het me ei o -Depreciation, salvage values, net operating working capital requilrements, and tax effects are all included in these cash flows. Both projects have 4-year lives, and they have risk characteristics simllar to the firms average project. Bellngers WACC Is 9% 360 230 300 Project A Project B 600 300 280350750 What is Project Deltas IRR? Do not round Intermediate calculations. Round your answer to two decimal places.
Answer #1

Project Delta's cashflows = Project A - Project B

Year 0 = -1050 - (-1050) = 0

Year 1 = 600 - 300 = 300

Year 2 = 360 - 280 = 80

Year 3 = 230 - 350 = -120

Year 4 = 300 - 750 = -450

IRR is the rate at which NPV = 0

0 = 300/(1+IRR)^1 + 80/(1+IRR)^2 - 120/(1+IRR)^3 - 450/(1+IRR)^4

By trail and error, IRR = 17.03%